OpenStudy (anonymous):

Find the amount that results from the investment. $10 invested at 6% compounded continuously after a period of 2 years. Plz I need help with this, I would like the steps..

OpenStudy (anonymous):

(10(0.6)24)+10=T (P(R)t)+P=T P=Principal R=Rate t=Time T=Total That is, if by "continuously" you mean monthly. If annually, then 't' = 2